OpenStudy (anonymous):

Solve, 2costheta-1=0, if 0

OpenStudy (anonymous):

2 cos theta= -1 =1/2 theta = 60 or theta = 300 degrees can someone verify, thx.

OpenStudy (ranga):

Is your first line supposed to be 2 cos theta= +1 (instead on minus 1?)

OpenStudy (anonymous):

no

OpenStudy (ranga):

First line in your solution.

OpenStudy (anonymous):

2 cos theta -1 = 0

OpenStudy (ranga):

theta = 60 or theta = 300 degrees is correct.

OpenStudy (anonymous):

thanks! you are brilliant

OpenStudy (anonymous):

2sin^2x+3sinx-2=0, if 0<x<2pi

OpenStudy (ranga):

You did all the work correctly in all three problems. I am just verifying what you have already done correctly. Kudos to you.

OpenStudy (anonymous):

I came up witth -pi/6 & 7pi/6 or -30 degrees and 210 degrees

OpenStudy (ranga):

I am getting sin(x) = 1/2 or x = 30 or 150 degrees

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

w x between 0 and 2 pi?

OpenStudy (anonymous):

ya you're right actually... i made a careless mistake

OpenStudy (anonymous):

(2sin(x)+1)(sin(x)-2=0

OpenStudy (ranga):

2sin^2x+3sinx-2=0 2sin^2x+4sinx - sinx -2=0 2sinx(sinx + 2) - 1(sinx+2) = 0 (2sinx-1)(sinx+2) = 0 sinx = 1/2 or sinx = -2 (which is not possible) So sinx = 1/2 x = 30 or 150 degrees and both fall in the domain 0<x<2pi

OpenStudy (anonymous):

so sin x=1/2 with x at 30 or 150

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

sin x-cos x=1, if 0<x<2pi i got x=pi/2 , 3pi/2 or x=pi

OpenStudy (ranga):

I have to go now. But you seem to know your stuff and so you will do fine.

OpenStudy (anonymous):

thanks

OpenStudy (ranga):

sin x-cos x=1, if 0<x<2pi x = pi/2 or x = pi The other solution you got x=3pi/2, if you put it back in the original equation sinx - cosx = 1 you will notice it does not satisfy the equation. It gives -1. The reason is because you probably squared both sides and so it will satisfy the squared equation but not the original equation and so you discard them.